What will be the expected fringe shift of earth in macheal moley exp. If the effective length of path is 8m and wave length of light used is 8000amstrung (v. Velocity of earth is 3×10^4
1
Expert's answer
2019-07-31T10:26:27-0400

The fringe shift of earth is given by formula:

"n = \frac{2×L×v ^2} {λ× c ^2} (1)"

In our case, L=8 m, v=3×104 m/s, λ=8000×10-10 m

Using (1) we get: n=0.2

Answer: 0.2